# Systematically Improvable Numerical Atomic Orbital Basis Using Contracted Truncated Spherical Waves

### 14 Mar 2026 v1

@@ III Results and Discussion
We first find the truncation radius $r_{\mathrm{c}}$ and the included maximal angular momentum ($l_{\max}$) of NSW for elements that can converge the total energy error with respect to PW $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ in dimer systems. Then we contract the NSW to build the NAO by minimizing the spillage defined in eqn 11, and compare the prediction errors of the properties with respect to PW and NSW (only for molecular systems). All our Density Functional Theory (DFT) calculations were performed with ABACUS (Atomic-orbital Based Ab-initio Computation at USTC) code [51, 54] 3.8.4. Perdew-Burke-Ernzerhof (PBE)  [76] functional was used for describing the exchange-correlation between electrons, and the interaction from ion cores and core electrons was represented with SG15 Optimized Norm-Conserving Vanderbilt (ONCV) pseudopotential [28, 86]. For isolated systems, enough vacuum was added in three directions (X/Y/Z) to avoid any overlap between occupied states of periodic images. Dipole correction was applied to those asymmetric molecules to eliminate dipole-dipole interactions between images. Additionally, for dioxygen and disulfur, we use constraint DFT (CDFT) functionality that explicitly constrains the difference in number of electrons between two channels (“spin-up” and “-down”) to 2, together with a spin-polarized KS-DFT calculation to obtain the most energy-favored states. For bulk systems, the $\mathbf{k}$-point sampling was done on a uniform grid (Monkhorst-Pack method [65]) with spacing $2\pi\times$0.08 Bohr-1, a Gaussian-type smearing on the electronic population used with a spread of 0.015 Ry to improve the Self-Consistent Field (SCF) convergence efficiency. For relaxation tasks, the Broyden–Fletcher–Goldfarb–Shanno (BFGS) and conjugate gradient (CG) algorithms are employed for the ion and ion-cell, respectively. The convergence thresholds of forces and stresses were set at 0.001 Ry/Bohr and 0.1 kbar.
@@ III.0.1 Converging truncated spherical wave parameters
With the given kinetic energy cutoff for spherical waves, the maximal angular momentum $l_{\max}$ (which means that the included angular momenta range from 0 to $l_{\max}$) and a uniform real-space cutoff radius $r_{\mathrm{c}}$ can together define a unique set of NSWs. By increasing $r_{\mathrm{c}}$ and $l_{\max}$ of NSWs, systematic converging behaviors towards PW are observed (Fig 2). At given $l_{\max}$, $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ decreases when $r_{\mathrm{c}}$ expands and shows converging behavior. Interestingly, this behavior varies substantially across elements. For alkali metals such as sodium (Na) shown in Fig 2(a), a strong dependence on the cutoff radius $r_{\mathrm{c}}$ is always shown; for elements of group IV-VII such as sulfur (S) and fluorine (F) in Fig 2(b) and (c), $l_{\max}$ is always more dominant than $r_{\mathrm{c}}$. Only when f ($l_{\max}$ = 3) or components with higher angular momentum are included, $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ can reach the level of chemical accuracy (1 kcal/mol). Such results are consistent with the variation in the electronegativity of elements. In this paper, we select orbital super-parameters $l_{\max}$ and $r_{\mathrm{c}}$ as small as possible for each element that converges $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ to 0.1 kcal/mol (4.2 meV). The values of $l_{\max}$ and $r_{\mathrm{c}}$, and the numbers of basis functions of the NAOs and NSWs of each element are shown in Table 1.
@@ III.0.2 Molecules
According to the variational theorem, the total energy is the primary criterion for quantitatively assessing basis-set completeness. We benchmarked our built NAOs on the 11 previously defined molecular systems to comprehensively assess their ability to describe wavefunctions across different chemical environments. The PW basis set is employed to approximate the CBS, whose kinetic energy cutoff is set to the value that can converge the total energy, pressure and electron-structure well; thus, the energy differences between NAO and PW roughly indicate the error due to basis-set incompleteness relative to CBS. We also employ primitive basis functions (NSW) in our benchmark, with NSW serving as a reference for assessing the loss of completeness via contraction and energy differences between NSW and PW, thereby indicating the rationality of the choices of truncation radius and maximal angular momentum.
@@
However, it is not only the number of basis functions that can impact the basis completeness, but the angular momentum of basis functions also matters. pVTZ and pVQZ= always have a very near number of basis functions, according to Table 1, but the performance always differs. In the $\mathrm{Br}_{2}$ case, the number of basis functions of pVTZ (3s3p2d1f) and pVQZ= (4s4p3d) cases are 29 and 31, respectively, and the difference in the energy errors is 0.063 eV/atom, reflecting the significant difference in the ability to expand the exact wavefunction. Such a difference can be explained from the basis optimization based on pVTZ- (3s3p2d), because there are already many radial functions to reproduce the wavefunction at some level; those basis functions leave the landscape of the s/p/d-type basis function optimization flat or highly non-convex, with no residual significant information in the reference states that can be extracted. Alternatively, adding a single f-type basis function may be preferable, enabling further significant reduction in spillage. As a result, pVTZ has spillage $2.189\times 10^{-4}$, while pVQZ= has $7.561\times 10^{-4}$. Extending the range of states included in spillage (by increasing $n$ in Eqn. 11) can also mitigate the ill-conditioning in the basis-function optimization tasks. More discussion will be provided in the last section and appendix. On the other hand, the NSW basis sets serve as the primitive basis sets and exhibit the smallest energy errors, reaching as low as 2.9 meV/atom on average. For $\mathrm{CO}$ and $\mathrm{S}_{2}$ molecules, the energy errors of NSW are 12 and 1 meV/atom, respectively, which are below the jY basis in Figure 1 in LRH’s work (about 30 and 7 meV/atom).
@@
In addition to testing whether our NAOs can accurately describe wavefunctions, structural property prediction is the most straightforward and particularly meaningful among molecular properties. We performed benchmarks on 11 molecular systems by optimizing atomic coordinates and measuring bond lengths, and compared the results with those obtained with PW basis sets (Table 3). Here, NSW is still used to evaluate the precision loss introduced by contraction and truncation separately.
Table 3: Bond lengths of $1$ molecules calculated with NAOs, NSW, PW basis sets and experimental data (in Å)
@@
Describing wavefunctions for both coordinated and isolated atomic states at near-precision levels is challenging. It stands as a golden rule in transferability tests - relevant not only for DFT functionals, but also for basis sets. The error, arising from differences in the ability to expand exact wavefunctions between these states, is the origin of the well-known drawback of atomic-centered orbital basis sets, the basis-set superposition error (BSSE). Because PW fully avoids BSSE, we benchmark our NAOs and NSW using the PW-calculated values as the reference.
@@
The occupation $f_{n\mathbf{k}}$-weighted Frobenius norm of state-wise eigenvalue $\varepsilon_{n\mathbf{k}}$ (energy) errors between two methods ($A$ and $B$) to evaluate the precision of electronic structure prediction, where the geometric average of occupation $\tilde{f}_{n\mathbf{k}}$ is defined by
@@
in which $f_{n\mathbf{k}}\left(\varepsilon_{\mathrm{f}}^{A},\sigma\right)$ and $f_{n\mathbf{k}}\left(\varepsilon_{\mathrm{f}}^{B},\sigma\right)$ are occupation numbers evaluated under a given smeared distribution. We use Gaussian smearing throughout this paper. $\varepsilon_{\mathrm{f}}^{A}$ and $\varepsilon_{\mathrm{f}}^{B}$ are Fermi levels that can be calculated based on present energy levels calculated by methods A and B, with a given smearing width ($\sigma$). In this test, $\eta$ is evaluated between NAOs, NSW, and PW, to quantify the precision loss from contraction, real-space truncation, and angular momentum selection. Superscripts are added to distinguish between $\eta$ values with respect to the PW and NSW basis sets ($\eta^{\mathrm{PW}}$, $\eta^{\mathrm{NSW}}$).
In Figure 3(d), the violin of NSW is already invisible. Quantitatively, in Table B.1, the $\eta$ of NSW basis sets is only 0.7 meV on average, which indicates that the NSW can predict the electronic structure nearly identical to PW. Among all molecules, NSW can predict the energy level of LiH with an error of less than 0.1 meV. The molecule with the highest $\eta$ is CO, with a value of 3.3 meV, but it remains negligible. For NAOs, the $\eta$ also shows a consistent decrease trend with respect to the increase of basis functions. Interestingly, although the errors of basis sets are almost negligible overall, and pVTZ is always the most out-performing NAOs except for $\mathrm{Li}_{2}$, $\mathrm{LiH}$ and $\mathrm{Na}_{2}$ molecules, where the pVQZ= basis sets have $\eta$ values only about half of the pVTZ basis sets. A closer examination of the difference of radial function profiles reveals that, pVQZ= always improves the pVTZ- basis set on the description of those more delocalized states, except for Li, where the newly added 7- and 8-th s-type radial functions are highly localized. Note that the pseudopotential of Li used in this work is of the semi-core type, where the 1s electrons are also considered as the “valence” electrons for transferability concerns. The improvement, then, can be rationalized as an improvement on the description of the 1s electrons.
@@ III.0.3 Bulks
In addition to static structural properties such as the lattice constant, a key mechanical property is the bulk modulus, which serves as the central quantity in solid equation-of-state (EOS) fitting and is closely related to the trace of the full stress tensor. Bulk modulus quantifies the rigidity to isotropic inflation-compression, defining structural behavior and mechanical response near volume equilibrium. Microscopically, the bulk modulus measures the local curvature of the bond-dissociation curves in the bulk system. Here, we benchmark the bulk moduli of 24 bulk samples using our built NAOs. The bulk modulus $B_{0}$ is calculated by fitting the Birch-Murnaghan equation [10] (in Eqn. 14), in which the volume $V$ ranges from 96% to 104% with a step size of 2% isotropically of the equilibrium volume $V_{0}$ obtained by a full relaxation with PW basis.
@@
We also introduce the $\Delta$-metric to quantify the difference between two EOS curves in a given volume range. With marks $A$ and $B$ distinguishing two EOS curves, the $\Delta\left(A,B\right)$ is defined as
@@
Among the $\Delta$ values of pVDZ, there are 35% (9 cases) below 1 meV/atoms, 62% (16 cases) below 1.5 meV/atoms. Such values are comparable with the 1 meV/atom threshold suggested by Prandini et al[81] in the benchmark of pseudopotentials against the all-electron, where pseudopotentials with $\Delta<$ 1 meV/atom are considered to be reliable. When the NAO increases to pVTZ-, 58% of cases have $\Delta$ values below the 1 meV/atom. For pVTZ, there are only the LiF, ZnO, ZnOw, and ZnS that have $\Delta$ values still larger than 1 meV/atom, while for ZnO, its value is already comparable with other atom-centered orbital-based DFT codes [12].
@@
As shown in Figure 4(f), all NAOs have average bandgap error very close to 0, and in Table 9, all NAOs have MAE not larger than 0.04 eV, which are near or below the standard deviation of the experimental bandgap measurement [61, 102]. From Table 9 and Figure 4(f), it can be seen that the introduction of additional basis functions with restricted angular momenta firstly eliminates the error of the GaN case, suppresses the positive tail of the pVDZ violin, and leaves the LiF still at the end of the negative tail. The similar lengths of the negative tails of pVTZ- and pVTZ indicates the higher angular momentum basis functions do not contribute significantly in improving the bandgap prediction, this is reasonable because the cases where the atomic orbitals with angular momenta $l+2$ (where $l$ is the largest angular momentum of electron in atomic configuration) the contributes significantly to states near the VBM or CBM are always scarcely seen. By leveling up the pVTZ- to pVQZ=, the error of LiF is reduced to 0.04 eV, and the negative tail of the violin is largely suppressed.
@@
GaN is a III/V direct bandgap semiconductor commonly used in blue light-emitting diodes. We checked the band structure over a wider energy range, corresponding to excitation wavelengths below 50 nm, thereby covering most of the optical spectrum. As shown in Figure 7 (a), it is similar to the NaCl case that only pVDZ overestimates the eigenvalues at some $\mathbf{k}$-point on the lowest conduction band. When the basis set increases to pVTZ-, there is no qualitative difference with PW up to 10 eV above the Fermi level. Inclusion of f-type orbital for N and g-type for Ga slightly lowers the bands in the range from 15 to 20 eV above the Fermi level, bringing a 0.006 eV decrease in $\eta_{10}$. The additional inclusion on the virtual states into spillage generates the pVTZ-2V and 3V, improves part of precision, as depicted in Figure 7 (d) and (e), for very high-lying states at about 20 eV above the Fermi level, on the path of kpoints $\Gamma-\mathrm{M}-\mathrm{K}-\Gamma$, pVTZ-2V and sometimes the pVTZ-3V can have energy closer to PW than pVTZ and pVQZ=, but the contrary result can also be observed on the path $\mathrm{A}-\mathrm{L}-\mathrm{H}-\mathrm{A}$, this explains the rise in $\eta_{10}$ from pVTZ to pVTZ-2V and pVTZ-3V.

### 09 Apr 2026 v2

@@ III Results and Discussion
We first find that the truncation radius $r_{\mathrm{c}}$ and the included maximal angular momentum ($l_{\max}$) of NSW for elements that can converge the total energy error with respect to PW $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ in dimer systems. Next, we contract the NSW basis to construct NAOs by minimizing the generalized spillage defined in Eq. 11, and compare the prediction errors of various properties relative to PW and NSW results for molecular systems only. All the DFT calculations were performed with ABACUS 3.8.4 (Atomic-orbital Based Ab-initio Computation at USTC) code [53, 56]. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [78] was used, and ion-core and core-electron interactions are represented using the SG15 Optimized Norm-Conserving Vanderbilt (ONCV) pseudopotential. [28, 88]. For isolated systems, enough vacuum was added in three directions ($X/Y/Z$) to avoid any overlap between occupied states of periodic images. Dipole correction was applied to those asymmetric molecules to eliminate dipole-dipole interactions between images.
Additionally, for dioxygen and disulfur, we performed spin-polarized Kohn–Sham DFT calculations combined with constrained DFT (CDFT), which explicitly constrains the electron number difference between spin-up and spin-down channels to 2, in order to obtain the energetically most favorable states. For bulk systems, the $\mathbf{k}$-point sampling was done on a uniform grid (Monkhorst-Pack method [67]) with spacing $2\pi\times$0.08 Bohr-1. A Gaussian-type smearing on the electronic population was used with a spread of 0.015 Ry to improve the Self-Consistent Field (SCF) convergence efficiency. For relaxation, the Broyden–Fletcher–Goldfarb–Shanno (BFGS) and conjugate gradient (CG) algorithms were employed for the ion and ion-cell, respectively. The convergence thresholds of forces and stresses were set at 0.001 Ry/Bohr and 0.1 kbar.
@@ III.0.1 Converging truncated spherical wave parameters
With a given kinetic energy cutoff for spherical waves, the maximum angular momentum $l_{\max}$ and a uniform real-space cutoff radius $r_{\mathrm{c}}$ together define a unique set of NSWs. The angular momentum components included cover the range from 0 to $l_{\max}$. By increasing $r_{\mathrm{c}}$ and $l_{\max}$ of NSWs, systematic converging behaviors towards PW are observed, as shown in Fig 2. At given $l_{\max}$, $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ decreases when $r_{\mathrm{c}}$ expands and shows converging behavior. Interestingly, this behavior varies substantially across elements. For alkali metals such as sodium (Na) shown in Fig 2(a), a strong dependence on the cutoff radius $r_{\mathrm{c}}$ is always shown; for elements of group IV-VII such as sulfur (S) and fluorine (F) in Fig 2(b) and (c), $l_{\max}$ is always more dominant than $r_{\mathrm{c}}$. Only when $f$ functions corresponding to $l_{\max}$=3 or components with higher angular momentum are included can $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ reach the chemical accuracy level of 1 kcal/mol. Such results are consistent with the variation in the electronegativity of elements. We select orbital parameters $l_{\max}$ and $r_{\mathrm{c}}$ as small as possible for each element that converges $\epsilon^{\mathrm{PW}}_{\mathrm{NSW}}$ to 0.1 kcal/mol (4.2 meV). The values of $l_{\max}$ and $r_{\mathrm{c}}$, and the numbers of basis functions of the NAOs and NSWs of each element are shown in Table 1.
@@ III.0.2 Molecules
According to the variational theorem, the total energy serves as the primary criterion for quantitatively evaluating basis-set completeness. We benchmarked the constructed NAOs on the 11 previously defined molecular systems to comprehensively evaluate their performance in describing wavefunctions under various chemical environments. The PW basis set is employed to approximate the CBS, whose kinetic energy cutoff is set to the value that can converge the total energy, pressure and electron-structure well; thus, the energy differences between NAO and PW roughly indicate the error due to basis-set incompleteness relative to CBS. We also employ primitive basis functions (NSW) in our benchmark, with NSW serving as a reference for assessing the loss of completeness via contraction and energy differences between NSW and PW, thereby indicating the rationality of the choices of truncation radius and maximal angular momentum.
@@
However, basis-set completeness depends not only on the number of basis functions but also on their angular momentum. The pVTZ and pVQZ= basis sets contain similar numbers of basis functions, according to Table 1, but the performance differs. For example, in the $\mathrm{Br}_{2}$ case, the number of basis functions of pVTZ (3s3p2d1f) and pVQZ= (4s4p3d) cases are 29 and 31, respectively. In addition, the energy difference is 0.063 eV/atom, reflecting the significant difference in the ability to expand the exact wavefunction. This can be explained from the basis optimization based on pVTZ- (3s3p2d), because sufficient radial functions are already present to reproduce the wavefunction to a given precision; those basis functions leave the landscape of the $s/p/d$-type basis function optimization flat or highly non-convex, as no significant residual information remains to be extracted from the reference states. Alternatively, adding a single $f$-type basis function may be preferable, enabling further significant reduction in spillage. As a result, pVTZ has spillage $2.189\times 10^{-4}$, while pVQZ= has $7.561\times 10^{-4}$. Extending the range of states included in spillage (by increasing $n$ in Eq. 11) can also mitigate the ill-conditioning in the basis-function optimization tasks. More discussion will be provided in the last section and appendix. On the other hand, the NSW basis sets serve as the primitive basis sets and exhibit the smallest energy errors, reaching as low as 2.9 meV/atom on average. For $\mathrm{CO}$ and $\mathrm{S}_{2}$ molecules, the energy errors of NSW are 12 and 1 meV/atom, respectively, which are below the jY basis in Figure 1 in LRH’s work (about 30 and 7 meV/atom).
@@
Besides verifying that our NAOs can accurately describe wavefunctions, structural property prediction is among the most straightforward and meaningful molecular properties to evaluate We benchmarked the 11 molecular systems by optimizing atomic coordinates and computing bond lengths, and compared the results with those from plane-wave basis sets (Table 3). Here, NSW is also used to separately evaluate the accuracy loss caused by contraction and truncation.
Table 3: Bond lengths of 11 molecules calculated with NAOs, NSW, PW basis sets and experimental data (in Å)
@@
Describing wavefunctions for both coordinated and isolated atomic states with near-quantitative precision is challenging. This represents a key criterion in transferability tests, applicable not only to DFT functionals but also to basis sets. Errors originating from differences in how well the exact wavefunction can be expanded between these states underlie the well-known limitation of atomic-centered orbital basis sets: the basis-set superposition error (BSSE). Since plane-wave calculations are entirely free of BSSE, we benchmark our NAOs and NSW against reference values obtained with PW basis sets.
@@
This metric can be understood as the occupation $f_{n\mathbf{k}}$-weighted Frobenius norm of state-wise eigenvalue $\varepsilon_{n\mathbf{k}}$ errors between two methods (AA and $B$), which is capable of evaluating the electronic structure prediction consistency. In eqn. 12, the minimization of $\eta$ by varying the level shift $\omega$ ensures the maximal alignment between two band structures, and the geometric average of occupation $\tilde{f}_{n\mathbf{k}}$ is defined by
@@
in which $f_{n\mathbf{k}}\left(\varepsilon_{\mathrm{f}}^{A},\sigma\right)$ and $f_{n\mathbf{k}}\left(\varepsilon_{\mathrm{f}}^{B},\sigma\right)$ are occupation numbers evaluated under a given smeared distribution (Gaussian smearing is used throughout this paper), $\varepsilon_{\mathrm{f}}$ and $\sigma$ denote the Fermi level and smearing width, respectively. In this test, $\eta$ is evaluated between NAOs, NSW, and PW, to quantify the precision loss from contraction, real-space truncation, and angular momentum selection. Superscripts are added to distinguish between $\eta$ values with respect to the PW and NSW basis sets ($\eta^{\mathrm{PW}}$, $\eta^{\mathrm{NSW}}$).
In Figure 3(d), the violin of NSW is invisible. Quantitatively, in Table B.1, the $\eta$ of NSW basis sets is only 0.7 meV on average, which indicates that the NSW can predict the electronic structure nearly identical to PW. Among all molecules, NSW can predict the energy level of LiH with an error of less than 0.1 meV. The molecule with the highest $\eta$ is CO, with a value of 3.3 meV, but it remains negligible. For NAOs, the $\eta$ also shows a consistent decrease trend with respect to the increase of basis functions. Interestingly, although the errors of basis sets are almost negligible overall, and pVTZ is always the most out-performing NAOs except for $\mathrm{Li}_{2}$, $\mathrm{LiH}$ and $\mathrm{Na}_{2}$ molecules, where the pVQZ= basis sets have $\eta$ values only about half of the pVTZ basis sets. A closer examination of the difference of radial function profiles reveals that, pVQZ= always improves the pVTZ- basis set on the description of those more delocalized states, except for Li, where the newly added 7- and 8-th s-type radial functions are highly localized. Note that the pseudopotential of Li used in this work is of the semi-core type, where the 1s electrons are also considered as the “valence” electrons for transferability concerns. The improvement, then, can be rationalized as an improvement on the description of the 1s electrons.
@@ III.0.3 Bulks
In addition to static structural properties such as the lattice constant, a key mechanical property is the bulk modulus, which serves as the central quantity in solid equation-of-state (EOS) fitting and is closely related to the trace of the full stress tensor. Bulk modulus quantifies the rigidity to isotropic inflation-compression, defining structural behavior and mechanical response near volume equilibrium. Microscopically, the bulk modulus measures the local curvature of the bond-dissociation curves in the bulk system. Here, we benchmark the bulk moduli of 24 bulk samples using our built NAOs. The bulk modulus $B_{0}$ is calculated by fitting the Birch-Murnaghan equation [10] (in Eqn. 14), where $E$ is the energy, $V$ is the volume, $V_{0}$ is the equilibrium volume, and $B_{0}^{\prime}$ is the derivative of bulk modulus $B_{0}$ with respect to the pressure. In our tests, $V$ ranges from 96% to 104% with a step size of 2% isotropically of $V_{0}$ obtained by a full relaxation with PW basis.
@@
We also introduce the $\Delta$-metric to quantify the difference between two EOS curves in the volume range from 96% $V_{0}$ ($V_{\mathrm{m}}$) to 104% $V_{0}$ ($V_{\mathrm{M}}$). With marks AA and $B$ distinguishing two EOS curves, the $\Delta\left(A,B\right)$ is defined as
@@
Among the $\Delta$ values of pVDZ, there are 35% (9 cases) below 1 meV/atoms, 62% (16 cases) below 1.5 meV/atoms. Such values are comparable with the 1 meV/atom threshold suggested by Prandini et al[83] in the benchmark of pseudopotentials against the all-electron, where pseudopotentials with $\Delta<$ 1 meV/atom are considered to be reliable. When the NAO increases to pVTZ-, 58% of cases have $\Delta$ values below the 1 meV/atom. For pVTZ, there are only the LiF, ZnO, ZnOw, and ZnS that have $\Delta$ values still larger than 1 meV/atom, while for ZnO, its value is comparable with other atom-centered orbital-based DFT codes [12].
@@
As shown in Figure 4(f), all NAOs have average bandgap error close to 0, and in Table 9, all NAOs have MAE not larger than 0.04 eV, which are near or below the standard deviation of the experimental bandgap measurement [63, 104]. From Table 9 and Figure 4(f), it can be seen that the introduction of additional basis functions with restricted angular momenta firstly eliminates the error of the GaN case, suppresses the positive tail of the pVDZ violin, and leaves the LiF still at the end of the negative tail. The similar lengths of the negative tails of pVTZ- and pVTZ indicates the higher angular momentum basis functions do not contribute significantly in improving the bandgap prediction, this is reasonable because the cases where the atomic orbitals with angular momenta $l+2$ (where $l$ is the largest angular momentum of electron in atomic configuration) the contributes significantly to states near the VBM or CBM are always scarcely seen. By leveling up the pVTZ- to pVQZ=, the error of LiF is reduced to 0.04 eV, and the negative tail of the violin is largely suppressed.
@@
GaN is a III/V direct bandgap semiconductor commonly used in blue light-emitting diodes. We checked the band structure over a wider energy range, corresponding to excitation wavelengths below 50 nm, thereby covering most of the optical spectrum. As shown in Figure 7 (a), it is similar to the NaCl case that only pVDZ overestimates the eigenvalues at some $\mathbf{k}$-point on the lowest conduction band. When the basis set increases to pVTZ-, there is no qualitative difference with PW up to 10 eV above the Fermi level. Inclusion of f-type orbital for N and g-type for Ga slightly lowers the bands in the range from 15 to 20 eV above the Fermi level, bringing a 0.006 eV decrease in $\eta_{10}$. The additional inclusion on the virtual states into spillage generates the pVTZ-2V and 3V, improves part of precision, as depicted in Figure 7 (d) and (e), for high-lying states at about 20 eV above the Fermi level, on the path of kpoints $\Gamma-\mathrm{M}-\mathrm{K}-\Gamma$, pVTZ-2V and sometimes the pVTZ-3V can have energy closer to PW than pVTZ and pVQZ=, but the contrary result can also be observed on the path $\mathrm{A}-\mathrm{L}-\mathrm{H}-\mathrm{A}$, this explains the rise in $\eta_{10}$ from pVTZ to pVTZ-2V and pVTZ-3V.